d) Soil column resistance 9. rue o r False. A safe means of access/egress that requires no more n 25 feet of lateral travel must be provided for workers in trenches excavated deeper than 4 ft in depth. (5 points) 10. You are required to excavate to a depth of 14 feet to install a section of sewer main. Youve examined the subsurface soils and classify the upper 5 feet as Type C soil and the soil deeper than 5 feet as Type B soil. You are above the water.table and excavation dewatering will not be necessary. a) Sketch the typical section OSHA would permit you to use for this excavation that would minimize overall excavation limits. (5 points) 1 b) If the trench bottom is planned to be 4 feet wide, how wide will the top of this excavation be if excavated per this section? (5 points) Ft Page 5 of 7
Answer #1

OSHA specifies that the excavation slope be 1:1 for Type B soil and 1.5:1 for Type C soil. Here the Type B soil is 9 feet deep and Type C soil is 5 feet deep.

Hence the width required at the top of Type B soil (at 9 feet from bottom) is tan 45 = opposite side / adjacent side, that is 1 = 9 feet / adjacent side, which gives a width of 9 feet on both sides. Hence the width at top of soil B is 4 + 9 + 9 = 22 feet.

Width required at top of soil C (at the top of excavation) is tan 34 = opposite side / adjacent side, that is 0.67 = 5 feet / adjacent side, that gives adjacent side = 5/0.67 = 7.46 on each side.

Hence, excavation width at the top of excavation = 22 + 7.46 + 7.46 = 36.92 or 37 feet